### v1

# New Methods for Offline GstLAL Analyses

Abstract
In this work, we present new methods implemented in the GstLAL offline gravitational wave search. These include a technique to reuse the matched filtering data products from a GstLAL online analysis, which hugely reduces the time and computational resources required to obtain offline results; a technique to combine these results with a separate search for heavier black hole mergers, enabling detections from a larger set of gravitational wave sources; changes to the likelihood ratio which increases the sensitivity of the analysis; and two separate changes to the background estimation, allowing more precise significance estimation of gravitational wave candidates. Some of these methods increase the sensitivity of the analysis, whereas others correct previous mis-estimations of sensitivity by eliminating false positives. These methods have been adopted for GstLAL’s offline results during the fourth observing run of LIGO, Virgo, and KAGRA (O4). To test these new methods, we perform an offline analysis over one chunk of O3 data, lasting from May 12 19:36:42 UTC 2019 to May 21 14:45:08 UTC 2019, and compare it with previous GstLAL results over the same period of time. We show that cumulatively these methods afford around a 50% - 100% increase in sensitivity in the highest mass space, while simultaneously increasing the reliability of results, and making them more reusable and computationally cheaper.
I Introduction
By detecting gravitational waves from the merger of compact objects like black holes and neutron stars, the LIGO Scientific, Virgo and KAGRA Collaborations (LVK) have revolutionized the field of gravitational wave astronomy. GW150914 was the first such detection of a GW Abbott et al. (2016a), and since then close to 250 GWs have been detected Collaboration (2024). While individual detections can often yield important scientific results, e.g., GW150914 Abbott et al. (2016b), GW170817 Abbott et al. (2018, 2019a), GW200105 and GW200115 Abbott et al. (2021a), even richer scientific implication can be extracted from a collection of GW detections. To this end, the LVK maintains a catalog of transient GW detections, called the Gravitational Wave Transient Catalog (GWTC). To date, four versions of the catalog have been released: GWTC-1 Abbott et al. (2019b), GWTC-2 gwt (2021), GWTC-2.1 The LIGO Scientific Collaboration et al. (2021), and GWTC-3 Abbott et al. (2021b). These help inform results related to testing the theory of general relativity Abbott et al. (2019c); Yang et al. (2020); Abbott et al. (2021c); Ezquiaga (2021), cosmology 245 and 247 (2017); Collaboration and the Virgo Collaboration (2021a), black hole properties and formation mechnisms Wu et al. (2020); Wang et al. (2021); Bouffanais et al. (2021), and binary population and merger rates Collaboration and the Virgo Collaboration (2019, 2021b); Abbott et al. (2023).
These results are enabled by GW search pipelines, by detecting GWs in the data produced by GW detectors, such as the two LIGO detectors Aasi et al. (2015), the Virgo detector Acernese et al. (2015), and the KAGRA detector Akutsu et al. (2020). The contents of the catalog are usually compiled from rigorous high-latency “offline” GW searches rather than low-latency “online” ones. GstLAL Messick et al. (2017); Sachdev et al. (2019); Cannon et al. (2020); Hanna et al. (2020) is one such GW search pipeline. Like other modeled GW searches, it makes use of a bank of waveform templates predicted by general relativity. These templates are cross-correlated against the data in a process called matched filtering to calculate the signal-to-noise ratio (SNR). GstLAL performs matched filtering in the time domain Messick et al. (2017); Sachdev et al. (2019); Cannon et al. (2012) This process is used to identify periods of time possibly containing GW signals, which are called “triggers”. Matched filtering is also used to inform the background data, against which triggers are ranked to evaluate their significance. A likelihood ratio (LR) Cannon et al. (2015); Tsukada et al. (2023) is calculated as the ranking statistic. Triggers with a high LR are called GW candidates. The LRs of candidates are then compared with the LR statistics of background triggers, and using the livetime of the analysis, a false alarm rate (FAR) is calculated for every candidate. Some of these techniques are used by other GW search pipelines, such as IAS Venumadhav et al. (2019); Zackay et al. (2021), MBTA Aubin et al. (2021); Adams et al. (2016), PyCBC Canton et al. (2021); Davies et al. (2020); Nitz (2018), and SPIIR Chu (2022, 2017).
This paper is structured as follows. In Sec. II, we describe the GstLAL offline workflow. Sec. III is dedicated to describing the new features introduced in the GstLAL offline workflow in the lead up to, and during the fourth observing run (O4) of the LVK. Sec. IV contains the results of the tests performed to evaluate these features, and to compare them against third observing run (O3) results.
II Overview of the GstLAL Offline Analysis
The GstLAL offline analysis can be broadly divided into two stages, each of which is further divided into smaller stages. These are:
1.
setup stage
template bank creation stage
power spectral density (PSD) measurement stage
singular value decomposition (SVD) of templates and template whitening stage
2.
data processing stage
matched filtering stage
rank stage
II.1 Template bank creation stage
For O4, two GstLAL template banks, namely the stellar-mass black hole and intermediate-mass black hole (IMBH) banks were generated using manifold (Hanna et al., 2023; Hanna, 2024), which is a binary-tree approach to template bank generation. All templates in both banks neglect eccentricity, higher-order modes, precession and matter effects. They additionally assume that the spins of each component object are aligned with an orbital angular momentum of the binary. This is done to limit the dimensionality of the bank, and hence the number of templates contained in it. While both of these banks are used to implement all-sky searches, in order to be consistent with internal conventions, in this work we will refer to the former search as the “AllSky search”, while the latter will continue to be called the IMBH search.
II.1.1 AllSky template bank
The AllSky template bank includes templates in the binary neutron star (BNS), neutron star–black hole binary (NSBH), and binary black hole (BBH) parameter spaces (Ewing et al., 2023), consisting of templates in total. This bank is used for the GstLAL online analysis, as well as for the bulk of the GstLAL offline results. It extends from component masses of to . The absolute values of dimensionless spin components are capped at 0.99 for components of the binary template that are black holes (defined as having a mass above ), and 0.05 for components that are neutron stars (defined as having a mass below ). Specifications of the AllSky template bank are shown in Tab. 1, and Fig. 1 shows a representation of the AllSky bank in the - space. Readers are referred to Ref. Sakon et al. (2022) for more details about the AllSky template bank.
Parameter | AllSky Template Bank |
Primary mass, | |
Secondary Mass, | |
Mass ratio, | |
Total mass, | |
Dimensionless spin, , for | |
Dimensionless spin, , for | |
Lower frequency cut-off | 15 Hz 111Due to enforcement of a maximum duration of 128 seconds, some lower mass tempaltes have a higher lower frequency cut-off |
Higher frequency cut-off | 1024 Hz |
Waveform approximant | IMRPhenomD |
Minimum match | 97 |
PSD | O4 projected PSD 222lig (2022) |
Total number of templates | 1815963 |
In comparison, the GstLAL O3 AllSky template bank covered component masses . It also limited templates to spin-aligned systems and set an upper limit of on the magnitude of the spin for component masses below . The spin magnitudes for component masses above were set to . The O3 template bank used a minimum match value of 99% for the BNS space, in contrast to the O4 value of 97%. The parameter choices of the O3 template bank is shown in Table II of Ref. gwt (2021). As a result of different template bank generation algorithms and parameter choices, the O3 AllSky bank consisted of templates.
The upper mass limit of the O4 AllSky bank is lower as compared to that of the O3 one. This was intentionally done, with the O4 IMBH bank (explained below) covering this space instead. The other differences in the O4 AllSky bank as compared to the O3 one were motivated by better template coverage of parameter spaces of interest with the new manifold method, while still minimizing the number of templates required.
II.1.2 IMBH template bank
In contrast to the AllSky bank, the IMBH bank is much smaller, and is only used to augment the GstLAL offline results in the IMBH parameter space. The IMBH bank covers the mass parameter space higher than what the AllSky template bank targets, as shown in Table 2. The lower limit of is set to such that the IMBH bank is an extention to the AllSky bank and the templates of the two template banks do not overlap in the - space, as shown in Fig. 1. The dimensionless spins extend from -0.7 to 0.99. The reason for the lower cutoff (-0.7) to be higher than that of the AllSky bank (-0.99) is to prevent the ringdown frequency of the templates from getting close to the lower frequency cutoff for matched filtering used by the IMBH search (10 Hz). Details of the parameters of the IMBH bank are shown in Tab. 2. Fig. 1 also shows a plot of the IMBH bank in the - space.
Parameter | IMBH Template Bank |
Primary mass, | |
Secondary Mass, | |
Mass ratio, | |
Total mass, | |
Dimensionless spin, | |
Lower frequency cut-off | 10 Hz |
Higher frequency cut-off | 1024 Hz |
Waveform approximant | IMRPhenomD |
Minimum match | 99 |
PSD | O4 projected PSD 333lig (2022) |
Total number of templates | 14728 |
II.1.3 Population model
Population models provide weights to templates which represent our prior knowledge of the astrophysical distribution of GW sources (Fong, 2018). This is used in the likelihood ratio (LR) calculation, as well as to compute probabilities of astrophysical origin of candidates (Ray et al., 2023). In O4, the population models for both the AllSky and IMBH banks were generated using manifold. They implement a Salpeter mass function to assign weights to templates, defined as:
For both the AllSky and IMBH population models, the value of is set to 0.8, which is slightly less than the minimum mass in the combined AllSky+IMBH template bank. The Salpeter function was chosen as a simple equation that works well over a large range of masses to approximate the astrophysical distibution of GW sources inferred by previous studies Collaboration and the Virgo Collaboration (2019, 2021b); Abbott et al. (2023).
II.2 PSD measurement stage
The PSD is the frequency representation of detector noise, and is used in the matched filter calculation. This is done by whitening both the data and the templates with the PSD measured from the data. GstLAL has the ability to measure the PSD of the data in real time and dynamically whiten the data during the matched filtering stage. However, this ability does not extend to template whitening, and the templates need to be whitened before the analysis starts. The PSD used for this purpose is measured from the data in the PSD measurement stage. The data is first divided into smaller pieces of up to 8 hours depending on the continuity of segments when each detector is operational. The PSD over each such piece of data is individually measured, and the median over each frequency bin is taken, which removes any transient noise features contributing to the PSD of a single piece of data. The resultant PSD is used to whiten the templates. More details about PSD measurement can be found in Ref. Messick et al. (2017).
[…]
Next, as discussed in Ref. Sakon et al. (2022), the template bank is sorted by linear combinations of their Post-Newtonian phase coefficients Morisaki and Raymond (2020) and split into “template bins”, each of around 1000 templates, which are whitened using the median PSD described above. These template bins further go through a process called SVD to save computational cost of the matched filtering Cannon et al. (2010). The templates in one such template bin are processed together for the purpose of matched filtering and background estimation, constituting a single matched-filtering job, and hence this template binning allows for a large-scale parallelization in both online and offline workflows.
II.4 Matched filtering stage
Each template within a template bin is matched filtered with the data, producing triggers. Alongside SNR, a signal consistency statistic, called is also calculated for every trigger Messick et al. (2017). Triggers with contribution from only one detector (referred to as a “single detector trigger”), during a time when multiple detectors were operational (referred to as “coincident time”), are considered to originate from noise Joshi et al. (2023). This is because GW signals are expected to be correlated across detectors, unlike noise. The SNR and statistics of noise triggers are added to the template bin’s common noise background, which is later one of the inputs used to calculate the LR for triggers.
In parallel, the GW candidates are extracted from the triggers. This is done by retaining only those triggers within a template bin that have the highest SNR in a 0.2 second window. A single GW signal can create triggers via multiple templates, and with current detector sensitivities, it is highly unlikely for multiple GW signals to occur in the same 0.2 second window /citeRelton2021. Hence, with this “clustering” over SNR, duplicate triggers from a single common signal are removed, reducing the amount of data downstream stages need to process. The processes of populating the background and generating candidates are done in parallel for every template bin.
II.5 Rank stage
In this stage, the significance of candidates is evaluated, and the results of the search are produced. First, the LR is calculated for every candidate forwarded by the matched filtering stage. One of the ingredients in calculating the LR is the SNR- background populated by the matched filtering stage. To model a smooth distribution of SNR- statistics, a kernel density estimate (KDE) is applied to the background, and the resulting probability density is used for LR evaluation. This gives us the probability that the (SNR, ) of a given candidate arises from noise. This is one of the many terms in the LR equation. More details on the LR calculation can be found in Ref. Tsukada et al. (2023).
Next, the candidate list is pruned again by performing a second round of clustering. This time, it is performed across template bins. The candidate with the highest LR in an 8 second window is retained. Apart from eliminating candidates originating from noise, this ensures that a single GW signal will have at most one candidate originating from it in the entire analysis.
Finally, to convert the LR to a FAR, we need the LR distribution of noise triggers from every template bin Cannon et al. (2013, 2015). If we use the same set of noise triggers as was used to populate the SNR- background, the LR distribution will not be well defined at higher LRs due to lower number of statistics. Additionally, since these noise triggers were obtained during the livetime of the analysis, the lowest FAR would be bound by per livetime. To solve this problem, GstLAL extrapolates the set of noise triggers from every template bin by drawing samples of SNR and , and assigning random templates as well as arrival times and phases. This is a computationally cheap operation, and under the assumption that the template, time and phase distribution of noise triggers is uniform, this process can effectively extrapolate the LR distribution of noise triggers to high LRs, enabling the FAR calculation of even the most significant candidates.
However, this process does not work particularly well for low LRs. To fix this, we can make one more assumption, that at low LRs, even the candidates originate from noise. Consequently, the low LR distribution of candidates is the same as the low LR distribution of noise triggers, and the former can be used to inform the latter. The method of doing so is called the “extinction model”. Once the LR distribution of noise triggers is calculated, it can be combined with the livetime of the analysis to produce a FAR for every candidate.
III New Methods
III.1 Online Rank
During every observing run, the full data are analyzed in near-real time by the GstLAL low-latency or “online” analysis. The various stages of the online analysis, up to the matched filtering stage, are largely similar to those of the offline analysis. Consequently, during O4, a new method has been adopted to outsource the GstLAL offline analysis’ matched filtering stage to the GstLAL online analysis. This method is called an “online rank”. A description of the differences between the GstLAL online and offline analyses, as well as a detailed description of the online rank method is provided in Ref. Joshi et al. (2025).
As described in Ref. Joshi et al. (2025), the online rank method significantly reduces the computational cost and time needed for offline results. Matched filtering is the vast majority of the computational burden of a modeled GW search. By eliminating the need to repeat matched filtering a second time for the offline analysis, this method affords a 50% reduction in computational cost over the course of an observing run, and specifically a 95 - 99.8% reduction in time required to obtain offline results.
III.2 Dropped Data Refiltering
In order to make the online rank results even more reliable and sensitive, we can augment the inputs to the online rank with triggers and background data from times that the online analysis dropped. To do this, for every job in the online analysis, we take the list of time periods when each job was functional and producing output. We then subtract that from the list of times a traditional offline analysis would have analyzed, which is calculated by external tools as an extra step in the setup stage of the offline workflow. This leaves us with a list of times that a particular job has dropped. We calculate such a “dropped data segments” for every job. Then, we can set up a traditional offline analysis using these dropped data segments. The only unusual aspect of this procedure is that while setting up this “dropped data refiltering” analysis, each job has its own segments list rather than a global one for every job. An example of such a list of dropped data segments is shown in Fig. 2
By combining the online rank inputs with the results of the dropped data refiltering analysis, we can be sure the online rank produces offline results for exactly the same periods of time that the traditional offline analysis would have. The typical amount of dropped data for any job is around 5% of the total time covered by the offline segments Joshi et al. (2025).
A point to note is that while we can add data to the online rank using this method, and we also have the ability to remove data to the granularity of the snapshots recorded by the online analysis (4 hours, see Ref. Joshi et al. (2025)), we cannot remove arbitrarily small amounts of data from an online rank. This means that we cannot incorporate more refined data quality information, such as small vetoed periods of time that only become available after the online analysis has finished.
III.3 IMBH Analysis
The AllSky template bank was designed to be a general-purpose bank for an online analysis. To obtain more comprehensive offline results, we want to augment the results of the AllSky search with a search for IMBH mergers. The GstLAL O4 offline results are obtained by performing a search over the data using the IMBH template bank described in Sec. II, which is then combined with the AllSky search into a single set of results using the procedue described in the following subsection.
As shown in Ref. Hanna et al. (2020); Tsukada et al. (2023), the LR contains a term to test the consistency of signals accross the network of detectors, for observed parameters like SNR and the coalescence time and phase. These parameters can only take on certain possible physical values for a network of detectors and follow specific correlations that indicate whether a GW trigger is of astrophysical origin. The correlation among these observables and their joint distribution depend on trigger’s template, and hence in principle this signal consistency test should be modified for every template. However, for the purpose of pre-computation efficiency, we adopt a BNS template with component masses of 1.4 , which is shown to be effective throughout Allsky template bank. In contrast, the mass range of the IMBH template bank is so high that this empirical fact might not be true, and hence for the IMBH analysis we use a template with component masses of 60 .
Other differences in search settings between the AllSky and the IMBH searches include different lower frequency cutoffs during matched filtering. AllSky uses 15 Hz, but the IMBH search uses 10 Hz as these binaries merge at lower frequencies. The Allsky search uses the TaylorF2 Buonanno et al. (2009) approximant upto a chirp mass of 1.73 for cost effectivity and the SEOBNRv4ROM Bohé et al. (2017) beyond that, while the IMBH search only uses the SEOBNRv4ROM as an approximant. Note that template placement for both banks was done with the IMRPhenomD approximant Khan et al. (2016), but we do not expect this inconsistency to be problematic. Similarly, the autocorrelation length used by the Allsky search for calculating the signal-based veto Messick et al. (2017) is 701 sample points upto a chirp mass of 1.73, and 351 above that. The IMBH search uses a value of 351 sample points throughout, due to a shorter duration of the waveforms. For this reason, the minimum number of samples included in a given time slice Messick et al. (2017) is 512 for IMBH and 2048 for Allsky. For similar reasons, while splitting the IMBH bank into template bins, the templates are sorted by their duration which is found to be more effective for this part of the parameter space, rather than sorting by linear combinations of their Post-Newtonian phase coefficients.
It is known that the candidates in the IMBH search are overwhelmingly single-detector candidates. This was true even for triggers, before any clustering was done. For example, in the test described in Sec. IV, out of the total candidates reported by the IMBH search, 92.75% were found to be single-detector candidates, and out of the set of triggers, 98.57% were single-detector triggers. The corresponding numbers for the AllSky search are 3.37% and 35.07% respectively. While the root cause of this is not well understood, the vastly different percentages of single-detector triggers in the AllSky and IMBH searches indicate that this arises from the matched filtering and coincidence formation of IMBH templates, rather than anything to do with the LR.
The LRs of single-detector candidates are notoriously difficult to estimate Callister et al. (2017), and so is the process of LR extrapolation of noise triggers. Consequently, the FARs of single-detector candidates are less reliable than those of multi-detector candidates, and a search with such a high percentage of single-detector candidates is unlikely to be functional. As a result, for GstLAL’s O4 offline results, the IMBH search ignores single detector triggers, and only processes those with multiple contributing detectors. The effects on the LR statistics of candidates if single-detector triggers are allowed to be processed are shown in Fig. 3.
III.4 Combining the AllSky and IMBH analyses
In order to combine Allsky and IMBH searches into a single search, one can naively treat every template bin from both the AllSky and IMBH searches as though they are part of a bigger AllSky+IMBH joint template bank. This means after each template bin assigns LRs to its candidates, these candidates would be clustered together based on their LR, and the noise LR statistics of each template bin will be combined via the extinction model. An implicit assumption of this method is that each template bin’s candidate and noise LR statistics in the noise regime are approximately the same. However, with the different parameter space and settings of the IMBH search as compared to AllSky, this assumption is no longer valid, and hence the naively combined results can be severely biased. To this end, we have developed a new scheme to combine the AllSky and IMBH analyses into a single search.
We first calculate a weight for each individual analysis, based on each analysis’ contribution to the final clustered set of candidates. This was done using injection campaigns for both analyses. Injections are simulated GW signals inserted into the data. This weight represents the number of GW signals we expect each analysis to contribute independent of the other, and is a combination of the analysis’ sensitivity, and the expected number of GW signals in each parameter space. The weights for the AllSky and IMBH analyses were found to be 0.94 and 0.06. This result was also verified by summing over the population model weights for the AllSky and IMBH templates after taking into consideration the different SNR detection thresholds of the two analyses.
Next, the FARs of each analysis’ candidates are scaled up by the inverse of their respective weights. Note that FARs are assigned to each analysis separately, and hence the FARs of candidates in the noise regime in both analyses are guaranteed to be distributed similarly as they only depend on the livetime of the search. Since the weights sum up to 1, this new set of candidates is guaranteed to have a consistent distribution of FARs in the noise regime, i.e., there will be on average livetime (expressed in hours) number of candidates with a FAR of one per hour or lower, and so on. Candidates from both analyses are then clustered together by retaining the lowest FAR candidate in an 8 second window. The LRs of the candidates are then re-calculated based on their new FARs, using the inverse of the mapping used by the AllSky analysis to convert LRs into FARs. In principle, any such mapping would have worked, and the AllSky mapping is chosen only for convenience.
III.5 Modularity and reusability of results
A general theme of the development work prior to and during O4 has been about reducing the duplication of computations done elsewhere. Apart from the online rank described in Sec. III.1, we have made the GstLAL offline workflow more modular and flexible. In general, offline analyses over large periods of data (such as GstLAL’s offline results for O4) are performed in small chunks, each of around 1 week of data. Alongside the ability to extract any amount of matched filtering data from the online analysis as specified by the user (to the granularity of 4 hours) Joshi et al. (2025), the GstLAL offline workflow also has the ability to use the data from any set of offline chunks specified by the user. Consequently, one can specify any period of data covered by any combination of online and offline analyses using any template banks, and they can be combined into a single result. Therefore, once a period of data has been matched filtered by a particular template bank via either an online (complemented by the dropped data refiltering) or offline analysis, it never needs to be filtered using that template bank again. This has greatly reduced the computational and time burden of GstLAL operations as well as development work during O4.
III.6 Ranking statistic improvements
GstLAL has adopted the likelihood ratio as the ranking statistic to evaluate the statistical significance of GW candidates Cannon et al. (2015); Tsukada et al. (2023). The likelihood ratio in GstLAL takes the form of
which represents the probability of obtaining a set of observable parameters, e.g., SNR for each detector , under the signal hypothesis () relative to that under the noise hypothesis (). Here, refers to the SNR of a single detector, and we will use this terminology for the rest of this subsection to maintain compatibility with others papers on this subject. The improvements in this ranking statistic prior to O4’s start are thoroughly described in Ref. Tsukada et al. (2023), which include the upgraded signal model and removal of signal contamination. In particular, the upgraded signal model achieved increase in the search sensitivity, and hence has led to a major contribution of GstLAL to LVK’s GW detections in O4 so far.
Apart from the improvements mentioned above, we have modified the noise model to make more accurate estimate of event’s significance in some part of () parameter space. As described in Ref. Tsukada et al. (2023); Joshi et al. (2023), GstLAL collects single-detector triggers during coincident time into the 2D histogram and estimates its noise model. Since templates of lighter binary systems such as BNS tend to not couple with non-Gaussian noise, e.g., glitches, most of noise triggers associated with such low-mass templates are expected to be modeled by the Gaussian noise component. For each detector, trigger’s SNR is calculated as square root of a quadrature sum of SNR values given by two orthogonal templates through matched-filtering, each of which follows the normal distribution. Therefore, the SNR-squared follows a chi-square distribution with two degree-of-freedom regardless of the observed value.
Also, according to the formalism of the upgraded signal model described in Ref. Tsukada et al. (2023), the expected value of in the noise model can be also given by substituting the mismatch factor , i.e., maximum mismatch between a hypothetical signal and template, into Eq.(40) of Ref. Tsukada et al. (2023),
where is a auto-correlation length used for calculation and is an auto-correlation function of a given template normalized so that . This approximation holds because an auto-correlation of compact binary coalescence (CBC) signals in general sharply peaks at the center where two templates exactly align, and hence . Therefore, triggers from the Gaussian noise component are distributed near and decaying toward higher SNR. Consistently, this behavior is illustrated in Fig. 5, which shows a noise model for one of the low-mass template bins collected for LIGO Hanford detector during the mock data campaign described in Ref. Chaudhary et al. (2023) with the lightblue dashed line as a contour. In contrast, a tail of the distribution, which extends toward higher SNR and , is known as non-Gaussian noise component.
However, during the online analysis of O4, we discovered that the Gaussian component of the noise model deviates from the expected chi-square distribution significantly at because very few or zero noise trigger populates such a higher SNR regime and the KDE applied as a smoothing process does not complement the lack of triggers sufficiently. See Fig. 5 where a 2D noise model is sliced at the contour and projected onto the dimension. Note that the expected chi-squared distribution shown as a black curve is given by
Although, this disagreement can eventually improve if we keep the analysis running and collect noise triggers long enough, early phase of the online analysis is likely to overestimate an event’s significance due to this bias and send out a public alert, e.g., S240422ed Ligo Scientific Collaboration et al. (2024a)444The statistical significance of this candidate was subsequently estimated to be lower after collecting more noise triggers and the update alert was sent Ligo Scientific Collaboration et al. (2024b). Also, we cannot guarantee that this bias does not exist even in an offline analysis where noise triggers collected entirely from single (or subset of) observing run are considered. Therefore, the objective here is to modify the tail of the Gaussian noise component such that it can produce more accurate noise model even with fewer collected triggers and eventually prevent a potential false claim of GW detections.
To this end, after collecting enough noise triggers, we extrapolate the trigger counts along a given iso- contour using the bulk of the Gaussian component so that its tail part follows the expected distribution shown in Eq. (4). This process is iterated over a range of and finally the KDE is applied to produce a smoother distribution. The green curve in Fig. 5 shows that the noise model after this extrapolation closely follows the expected distribution in black curve. Although it can be slightly above the black curve, one can see that it still provides much more accurate estimate of the event’s significance than the red curve without the extrapolation at .
III.7 New Extinction Model
A set of candidates undergo two rounds of clustering, the first with other candidates within the template bin, and the second with candidates from all template bins. These clustering processes alter the LR distribution of candidates. Consequently, the LR distribution of the noise triggers which gets used to convert the LR of candidates to FARs also needs to undergo these clustering processes. The new extinction model quantifies the change in LR distribution due to clustering, and applies it to the LR distribution on noise triggers. To mimic clustering, this is performed in two steps, the first within the noise triggers of a template bin, and the second with noise triggers across template bins. Additionally, the new extinction model enforces that any template bin’s contribution to the noise LR distribution is proportional to its contribution to the LR distribution of candidates. Both of these features ensure that the FARs assigned to candidates are more accurate than the old method, i.e., stitching the LR distribution of candidates below some LR threshold to the LR distribution of noise triggers Messick (2019).
Let be the histogram of noise triggers that GstLAL stores in order to estimate the complementary cumulative distribution function of noise LRs, . The two are related in the following way:
where A is some normalization constant, and the subscript denotes the clustered version of the respective functions. Equivalently,
The new extinction model assumes that the process of obtaining triggers above a certain LR threshold in the clustering interval is a Poisson process, with Poisson rate given by:
where is another normalization constant. Hence the probability that some trigger with LR = survives clustering, i.e. that there are no triggers with a higher LR than in the clustering interval is the Poisson probability for zero events.
Hence, we can add the effect of clustering to , by multiplying it with the survival probability.
Substituting Eq. (10) in Eq. (6), it follows
where = 0. Since GstLAL deals with the complementary cumulative histogram of LRs, , rather than probability density, we can convert this to the histogram form by adding another normalization constant. After absorbing relevant constants in and , we get:
In order to find the constants and , the new extinction model performs a curve fit of to the LR histogram of candidates. This is first done within a template bin, and after all of those clustered bin-specific noise histograms are added together, it is performed across template bins. The constant takes care of the relative contributions of template bins to the candidate set, whereas the constant takes care of the effect of clustering on the set of candidates. The curve fitting is performed from the 50th percentile to the 99th percentile of the complementary cumulative histogram of candidate LRs. This is empirically known to be a region of well-modeled noise candidates. The uncertainties in the curve fitting can be calculated from the assumption that a particular realization of the value of is Poisson distributed. Hence,
The effect of applying the new extinction model on the noise triggers is shown in Fig. 6. We can see that in the noise regime (i.e. 50th percentile to the 99th percentile), the noise LR histogram with the new extinction model is very close to the candiate histogram. This shows that the new extinction model is successful in applying the effects of candidate clustering and relative contributions of template bins to the candidate set. By adopting the new extinction model, we get more accurate FARs, since the noise LR distribution used to assign FARs is more accurate than the old method. The new extinction model has also been adopted for GstLAL’s online operations in the second half of the fourth observing run (O4b).
IV Results
IV.1 Data set
To test the new methods described in Sec. III, and to serve as a trial run for GstLAL’s O4 offline results, we set up a GstLAL analysis over 1 chunk of LIGO and Virgo O3 data. The data extends for approximately one week, from May 12 19:36:42 UTC 2019 to May 21 14:45:08 UTC 2019. Since Ref. Joshi et al. (2025) already tests the online rank feature and concludes it is equivalent to a traditional offline analysis, we do not test that feature here, and directly set up an offline analysis over the data instead of running an online analysis and setting up an online rank based on that. The offline analysis included a search with the AllSky bank, and a search with the IMBH bank, which were then combined into a single search using the procedure described in Sec. III, referred to as the “O4 analysis” hereafter. It also includes the LR improvements, the extrapolation in noise model and new extinction model. We then compared the results to GstLAL’s O3 offline results over the same period of data, referred to as the “O3 analysis” hereafter. Also, note that the O3 analysis uses the O3 template bank described in Sec. II.1.1. Both analyses also included the same injection campaign. The details of the distribution of injections in this campaign can be found in Ref. gwt (2021); Collaboration and Collaboration .
IV.2 Candidate lists
This set of data contains 6 GW candidates previously reported in Ref. gwt (2021); The LIGO Scientific Collaboration et al. (2021). Both the O4 and O3 analyses recover all 6 candidates in the list of top 10 candidates, as summarized in Tab. 3 and Tab. 4 respectively. The O4 search recovers all 6 confidently with a FAR below 1/month, whereas the O3 search only recovers 5 of the 6 confidently.
Rank | FAR (Hz) | SNR | UTC Time | Found Detectors | Operating Detectors | () | () | ||
1 | 24.51 | 2019-05-21 07:43:59 | H1,L1 | H1,L1 | 54.75 | 29.25 | -0.2011 | -0.2011 | |
2 | 13.35 | 2019-05-19 15:35:44 | H1,L1 | H1,L1 | 83.15 | 79.75 | 0.6033 | 0.6033 | |
3 | 14.14 | 2019-05-21 03:02:29 | H1,L1 | H1,L1,V1 | 213.1 | 137.4 | 0.8712 | 0.8712 | |
4 | 12.30 | 2019-05-13 20:54:28 | H1,L1,V1 | H1,L1,V1 | 50.36 | 48.29 | 0.6033 | 0.6033 | |
5 | 10.48 | 2019-05-17 05:51:01 | H1,L1 | H1,L1,V1 | 50.36 | 40.86 | 0.7889 | 0.7889 | |
6 | 10.88 | 2019-05-20 10:35:19 | L1,V1 | H1,L1,V1 | 442.1 | 218.2 | 0.8646 | 0.8646 | |
7 | 8.41 | 2019-05-14 06:54:16 | H1,L1 | H1,L1 | 54.75 | 57.08 | -0.2320 | -0.2320 | |
8 | 13.00 | 2019-05-15 06:27:00 | H1,L1 | H1,L1 | 543.3 | 63.58 | 0.2638 | 0.2638 | |
9 | 13.67 | 2019-05-14 04:01:29 | H1,L1 | H1,L1 | 414.9 | 144.7 | 0.8778 | 0.8778 | |
10 | 8.39 | 2019-05-18 20:43:46 | L1,V1 | H1,L1,V1 | 414.9 | 218.2 | 0.8646 | 0.8646 |
Rank | FAR (Hz) | SNR | UTC Time | Found Detectors | Operating Detectors | () | () | ||
1 | 24.43 | 2019-05-21 07:43:59 | H1,L1 | H1,L1 | 56.33 | 37.81 | -0.0015 | 0.4993 | |
2 | 11.68 | 2019-05-13 20:54:28 | H1,L1 | H1,L1 | 60.74 | 32.32 | 0.8530 | -0.1904 | |
3 | 13.03 | 2019-05-19 15:35:44 | H1,L1 | H1,L1 | 103.40 | 11.61 | 0.0925 | -0.9660 | |
4 | 10.36 | 2019-05-17 05:51:01 | H1,L1,V1 | H1,L1,V1 | 49.39 | 39.63 | 0.6712 | 0.8675 | |
5 | 12.94 | 2019-05-21 03:02:29 | H1,L1,V1 | H1,L1,V1 | 139.70 | 2.92 | -0.2471 | -0.0434 | |
6 | 8.42 | 2019-05-14 23:48:57 | L1,V1 | L1,V1 | 2.34 | 1.24 | -0.0130 | 0.0336 | |
7 | 8.93 | 2019-05-20 06:39:01 | H1,L1 | H1,L1 | 4.94 | 1.10 | 0.3486 | -0.0499 | |
8 | 8.33 | 2019-05-14 06:54:16 | H1,L1 | H1,L1 | 53.48 | 53.48 | -0.5578 | -0.5578 | |
9 | 9.63 | 2019-05-20 20:45:36 | H1,V1 | H1,L1,V1 | 21.42 | 18.44 | 0.6018 | 0.4931 | |
10 | 8.92 | 2019-05-17 14:43:16 | H1,L1 | H1,L1 | 19.96 | 1.24 | -0.9457 | 0.0449 |
IV.3 Sensitivity comparison
To compare the sensitivities of the O4 and O3 analyses, we use the sensitive volume-time, or VT of the two analyses as a metric. It represents the volume of 4-dimentional space-time where the search can typically identify GW signals. Since the times analyzed by both the O4 and O3 analyses is exactly the same, the VTs of the two analyses are a measure of their relative sensitivities. Furthermore, the VTs can be calculated for different mass ranges in order to estimate the relative sensitivities to different source classes. A plot of the ratio of the VT of the O4 analysis with and without the extrapolation in noise model, with the O3 analysis is shown in Fig. 7.
IV.4 Performance of individual features
IV.4.1 IMBH analysis
Here, we assess how much sensitivity we gain by combining the IMBH search with the AllSky search, compared to just the sensitivity of the AllSky search. The VT ratio of the O4 combined AllSky+IMBH search to that of the VT of the O4 AllSky search is shown in Fig. 8. In addition to the 6% - 7% sensitivity improvement in the IMBH space as shown in Fig. 8, by adding the IMBH search to the AllSky one, the combined analysis becomes sensitive to a new parameter space, and has the ability to recover GW signals arising from intermediate-mass black hole mergers, which the AllSky search did not have by itself. This source class is particularly interesting and by detecting GWs from such mergers, we enable new scientific results Abbott et al. (2022).
IV.4.2 Extrapolation in noise model
As mentioned previously, the extrapolation in noise model removes false positives in GW candidates. This implies that real GW signals can also be downranked accordingly and not be recovered as significantly as it would be without the extrapolation. To this end, we compare the s between with and without the extrapolation to assess the signal recovery performance. Fig. 9 shows the of the O4 analysis with the noise model extrapolation as a function of the recovered FAR for each category of injections, being relative to that for without the extrapolation. Note that the sensitivity can decrease up to 15% for two lowest mass categories of the injections, e.g., blue and orange curves. This is because, as mentioned above, the noise model for lower mass templates tends to be dominated by the Gaussian component and hence the extrapolation is likely to be more impactful. In contrast, for heavier BBH templates, the non-Gaussian noise component overwhelms the Gaussian one and makes small changes at its tail part negligible. We also emphasize that this decrease in the search sensitivity does not necessarily indicate the lowered performance overall, but rather that the signal recovery without the extrapolation is overestimated and risks potential false positives.
IV.4.3 Reliability of results
The new extinction model and LR improvements ensure the FARs of candidates are more accurate, and the extrapolation in noise model reduces the risk of false positives in the candidate set. All of these contribute to the O4 results being more reliable then before. This effect is difficult to measure since wrongly assigned FARs and false positives are rare. One possible metric is to see how many candidates in the injection campaign with a FAR below some threshold do not correspond to an injected GW signal. With a FAR threshold of one per day, this number for the O3 search is 53, representing 0.07% of all candidates. The corresponding number for the O4 search is 28, representing 0.03% of all candidates. For a higher FAR threshold of one per hour. the numbers for the O3 and O4 searches is 276 (0.35%) and 132 (0.15%). This shows that the new O4 methods are successful in lowering the number of false positives.
The O4 search includes the IMBH search. This parameter space is known to be particularly vulnerable to loud noise transients or non-gaussian transient “glitches” Ghosh et al. (2024); Macas et al. (2022). If we only look at the O4 AllSky search, for a more direct comparison with O3, the numbers for the one-per-day and one-per-hour FAR thresholds are 23 (0.03%), and 110 (0.13%), showing an even higher efficacy of removing false positives.
V Conclusion
In this work, we first gave an overview of the GstLAL offline analysis. This included desriptions of the GstLAL AllSky and IMBH template banks used in O4. We later described new methods in the GstLAL offline analysis introduced in the lead up to and during O4. These new methods are all used to obtain GstLAL’s offline results for O4.
The new methods include a way to outsource everything up to the matched filtering stage to an online analysis that had previously run on the same data. This, along with improvements to modularity of the workflow and reusability of results means that GW data only needs to be matched filtered with a given template bank only once during an observing run. The results of any such matched filtering, whether online or offline, and with any template bank can be combined into a single cohesive set of results.
The new methods also include a new IMBH offline search. This search only allows candidates with multiple contributing detectors, in order to make the search well-behaved and functional. A new method to combine the AllSky and IMBH searches was also introduced. It assigns a weight to each search by which the FARs of their candidates, calculated independently for each search, are scaled up. Afterwards, the set of candidates of the two searches can be combined in to single set of candidates.
Finally, we also discussed some changes to the LR calculation, introduced the extrapolation in noise model, as well as a new extinction model which enables a better estimation of the LR statistics of noise triggers.
To test these new methods, we set up an offline analysis over one week of O3 data, and compared the results with those obtained from the same data by GstLAL during O3. Both searches recovered the 6 previously reported GWs in the data in their list of top 10 candidates. The O4 search was able to do so confidently for all 6, whereas the O3 search managed it only for 5. We found that the sensitivity of the O4 search in the IMBH space increased by 50% - 100%. The VT in the BNS and NSBH space went down by 10% - 20% as a result of the extrapolation in noise model, which reduces false positives and downweighting a small number of real GW signals accordingly. Without the new KDE, those VTs are very close to the O3 ones.
The new methods described in this work variously increase the sensitivity, reliability, or reusability of the GstLAL offline results. As a result, we expect GstLAL’s offline results to significantly contribute to the LVK’s scientific results in O4.
Acknowledgements.
This research has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org/ ), a service of LIGO Laboratory, the LIGO Scientific Collaboration (LSC) and the Virgo Collaboration. We especially made heavy use of the LVK Algorithm Library. LIGO was constructed by the California Institute of Technology and the Massachusetts Institute of Technology with funding from the United States National Science Foundation (NSF) and operates under cooperative agreements PHYS- and PHY-. In addition, the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany supported the construction of Advanced Laser Interferometer Gravitational-Wave Observatory (aLIGO) and construction and operation of the GEO600 detector. Additional support for aLIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. The authors are grateful for computational resources provided by the LIGO Lab culster at the LIGO Laboratory and supported by PHY- and PHY, the Pennsylvania State University’s Institute for Computational and Data Sciences gravitational-wave cluster, and supported by OAC-, PHY-, PHY-, OAC-, OAC-, and PHY-. The authors thank Thomas Dent (and the LVK CBC AllSky group) for sharing initial investigations on the Gaussian noise component of the noise model discussed in Sec. III.6. The authors also thank Reed Essick (and the entire LVK Rates and Populations group) for developing the injection sets used in the calculation of the weights for the AllSky and IMBH analyses. LT acknowledges NASA 80NSSC23M0104 and the Nevada Center for Astrophysics for support. CH Acknowledges generous support from the Eberly College of Science, the Department of Physics, the Institute for Gravitation and the Cosmos, the Institute for Computational and Data Sciences, and the Freed Early Career Professorship.References
- Abbott et al. (2016a) B. P. Abbott et al. (LIGO Scientific, Virgo), Phys. Rev. Lett. 116, 061102 (2016a), arXiv:1602.03837 [gr-qc] . equation of state (EoS)
- Collaboration (2024) L. Collaboration, “Ligo-virgo-kagra cumulative detection plot - o1-o4b,” (2024).EoS
- Abbott et al. (2016b) B. P. Abbott et al. (LIGO Scientific and Virgo Collaborations), Phys. Rev. Lett. 116, 221101 (2016b).EoS
- Abbott et al. (2018) B. P. Abbott et al. (The LIGO Scientific Collaboration and the Virgo Collaboration), Phys. Rev. Lett. 121, 161101 (2018).EoS
- Abbott et al. (2019a) B. P. Abbott et al. (LIGO Scientific Collaboration and Virgo Collaboration), Phys. Rev. Lett. 123, 011102 (2019a).EoS
- Abbott et al. (2021a) R. Abbott et al. (LIGO Scientific, KAGRA, VIRGO), Astrophys. J. Lett. 915, L5 (2021a), arXiv:2106.15163 [astro-ph.HE] .EoS
- Abbott et al. (2019b) B. P. Abbott et al. (LIGO Scientific, Virgo), Phys. Rev. X 9, 031040 (2019b), arXiv:1811.12907 [astro-ph.HE] .EoS
- gwt (2021) Physical Review X 11, 021053 (2021).EoS
- The LIGO Scientific Collaboration et al. (2021) The LIGO Scientific Collaboration, The Virgo Collaboration, R. Abbott, et al., “Gwtc-2.1: Deep extended catalog of compact binary coalescences observed by ligo and virgo during the first half of the third observing run,” (2021).EoS
- Abbott et al. (2021b) R. Abbott et al. (LIGO Scientific, VIRGO, KAGRA), “GWTC-3: Compact Binary Coalescences Observed by LIGO and Virgo During the Second Part of the Third Observing Run,” (2021b), arXiv:2111.03606 [gr-qc] .EoS
- Abbott et al. (2019c) B. P. Abbott et al. (The LIGO Scientific Collaboration and the Virgo Collaboration), Phys. Rev. D 100, 104036 (2019c).EoS
- Yang et al. (2020) S.-C. Yang, W.-B. Han, and G. Wang, Monthly Notices of the Royal Astronomical Society: Letters 499, L53 (2020), https://academic.oup.com/mnrasl/article-pdf/499/1/L53/54638219/mnrasl_499_1_l53.pdf .EoS
- Abbott et al. (2021c) R. Abbott, H. Abe, F. Acernese, K. Ackley, N. Adhikari, R. Adhikari, V. Adkins, V. Adya, C. Affeldt, D. Agarwal, et al., arXiv preprint arXiv:2112.06861 (2021c).EoS
- Ezquiaga (2021) J. M. Ezquiaga, Physics Letters B 822, 136665 (2021).EoS
- 245 and 247 (2017) D. C. H. J. . K. V. . R. D. . T. L. . . S. D. . V. S. . Y. S. . . 245 and L. C. O. C. A. I. . . H. G. . . H. D. A. . . M. C. . . P. D. . V. S. . 247, Nature 551, 85 (2017).EoS
- Collaboration and the Virgo Collaboration (2021a) T. L. S. Collaboration and the Virgo Collaboration, The Astrophysical Journal 909, 218 (2021a).EoS
- Wu et al. (2020) S. Wu, Z. Cao, and Z.-H. Zhu, Monthly Notices of the Royal Astronomical Society 495, 466 (2020), https://academic.oup.com/mnras/article-pdf/495/1/466/33387647/staa1176.pdf .EoS
- Wang et al. (2021) H.-T. Wang, P.-C. Li, J.-L. Jiang, G.-W. Yuan, Y.-M. Hu, and Y.-Z. Fan, The European Physical Journal C 81, 769 (2021).EoS
- Bouffanais et al. (2021) Y. Bouffanais, M. Mapelli, F. Santoliquido, N. Giacobbo, U. N. Di Carlo, S. Rastello, M. C. Artale, and G. Iorio, Monthly Notices of the Royal Astronomical Society 507, 5224 (2021), https://academic.oup.com/mnras/article-pdf/507/4/5224/40392757/stab2438.pdf .EoS
- Collaboration and the Virgo Collaboration (2019) T. L. S. Collaboration and the Virgo Collaboration, The Astrophysical Journal Letters 882, L24 (2019).EoS
- Collaboration and the Virgo Collaboration (2021b) T. L. S. Collaboration and the Virgo Collaboration, The Astrophysical Journal Letters 913, L7 (2021b).EoS
- Abbott et al. (2023) R. Abbott et al. (KAGRA, VIRGO, LIGO Scientific), Phys. Rev. X 13, 011048 (2023), arXiv:2111.03634 [astro-ph.HE] .EoS
- Aasi et al. (2015) J. Aasi et al. (LIGO Scientific), Class. Quant. Grav. 32, 074001 (2015), arXiv:1411.4547 [gr-qc] .EoS
- Acernese et al. (2015) F. Acernese et al. (VIRGO), Class. Quant. Grav. 32, 024001 (2015), arXiv:1408.3978 [gr-qc] .EoS
- Akutsu et al. (2020) T. Akutsu et al., Progress of Theoretical and Experimental Physics 2021 (2020), 10.1093/ptep/ptaa125, 05A101, https://academic.oup.com/ptep/article-pdf/2021/5/05A101/37974994/ptaa125.pdf .EoS
- Messick et al. (2017) C. Messick et al., Phys. Rev. D 95, 042001 (2017), arXiv:1604.04324 [astro-ph.IM] .EoS
- Sachdev et al. (2019) S. Sachdev et al., “The GstLAL Search Analysis Methods for Compact Binary Mergers in Advanced LIGO’s Second and Advanced Virgo’s First Observing Runs,” (2019), arXiv:1901.08580 [gr-qc] .EoS
- Cannon et al. (2020) K. Cannon et al., “GstLAL: A software framework for gravitational wave discovery,” (2020), arXiv:2010.05082 [astro-ph.IM] .EoS
- Hanna et al. (2020) C. Hanna et al., Phys. Rev. D101, 022003 (2020), arXiv:1901.02227 [gr-qc] .EoS
- Cannon et al. (2012) K. Cannon et al., Astrophys. J. 748, 136 (2012), arXiv:1107.2665 [astro-ph.IM] .EoS
- Cannon et al. (2015) K. Cannon, C. Hanna, and J. Peoples, “Likelihood-Ratio Ranking Statistic for Compact Binary Coalescence Candidates with Rate Estimation,” (2015), arXiv:1504.04632 [astro-ph.IM] .EoS
- Tsukada et al. (2023) L. Tsukada, P. Joshi, S. Adhicary, R. George, A. Guimaraes, C. Hanna, R. Magee, A. Zimmerman, P. Baral, A. Baylor, K. Cannon, S. Caudill, B. Cousins, J. D. E. Creighton, B. Ewing, H. Fong, P. Godwin, R. Harada, Y.-J. Huang, R. Huxford, J. Kennington, S. Kuwahara, A. K. Y. Li, D. Meacher, C. Messick, S. Morisaki, D. Mukherjee, W. Niu, A. Pace, C. Posnansky, A. Ray, S. Sachdev, S. Sakon, D. Singh, R. Tapia, T. Tsutsui, K. Ueno, A. Viets, L. Wade, and M. Wade, (2023), arXiv:2305.06286 [astro-ph.IM] .EoS
- Venumadhav et al. (2019) T. Venumadhav, B. Zackay, J. Roulet, L. Dai, and M. Zaldarriaga, Phys. Rev. D 100, 023011 (2019).EoS
- Zackay et al. (2021) B. Zackay, L. Dai, T. Venumadhav, J. Roulet, and M. Zaldarriaga, Physical Review D 104 (2021), 10.1103/physrevd.104.063030.EoS
- Aubin et al. (2021) F. Aubin, F. Brighenti, R. Chierici, D. Estevez, G. Greco, G. M. Guidi, V. Juste, F. Marion, B. Mours, E. Nitoglia, O. Sauter, and V. Sordini, Classical and Quantum Gravity 38, 095004 (2021).EoS
- Adams et al. (2016) T. Adams, D. Buskulic, V. Germain, G. M. Guidi, F. Marion, M. Montani, B. Mours, F. Piergiovanni, and G. Wang, Classical and Quantum Gravity 33, 175012 (2016).EoS
- Canton et al. (2021) T. D. Canton, A. H. Nitz, B. Gadre, G. S. C. Davies, V. Villa-Ortega, T. Dent, I. Harry, and L. Xiao, The Astrophysical Journal 923, 254 (2021).EoS
- Davies et al. (2020) G. S. Davies, T. Dent, M. Tápai, I. Harry, C. McIsaac, and A. H. Nitz, Physical Review D 102 (2020), 10.1103/physrevd.102.022004.EoS
- Nitz (2018) A. H. Nitz, Physical Review D 98 (2018), 10.1103/PhysRevD.98.024050.EoS
- Chu (2022) Q. Chu, Physical Review D 105 (2022), 10.1103/PhysRevD.105.024023.EoS
- Chu (2017) Q. Chu, Low-latency detection and localization of gravitational waves from compact binary coalescences, Ph.D. thesis, The University of Western Australia (2017).EoS
- Hanna et al. (2023) C. Hanna, J. Kennington, S. Sakon, S. Privitera, M. Fernandez, J. Wang, C. Messick, A. Pace, K. Cannon, P. Joshi, R. Huxford, S. Caudill, C. Chan, B. Cousins, J. D. E. Creighton, B. Ewing, H. Fong, P. Godwin, R. Magee, D. Meacher, S. Morisaki, D. Mukherjee, H. Ohta, S. Sachdev, D. Singh, R. Tapia, L. Tsukada, D. Tsuna, T. Tsutsui, K. Ueno, A. Viets, L. Wade, and M. Wade, Phys. Rev. D 108, 042003 (2023).EoS
- Hanna (2024) C. Hanna, “manifold,” https://git.ligo.org/chad-hanna/manifold (2024).EoS
- Ewing et al. (2023) B. Ewing, R. Huxford, D. Singh, L. Tsukada, C. Hanna, Y.-J. Huang, P. Joshi, A. K. Y. Li, R. Magee, C. Messick, A. Pace, A. Ray, S. Sachdev, S. Sakon, R. Tapia, S. Adhicary, P. Baral, A. Baylor, K. Cannon, S. Caudill, S. S. Chaudhary, M. W. Coughlin, B. Cousins, J. D. E. Creighton, R. Essick, H. Fong, R. N. George, P. Godwin, R. Harada, J. Kennington, S. Kuwahara, D. Meacher, S. Morisaki, D. Mukherjee, W. Niu, C. Posnansky, A. Toivonen, T. Tsutsui, K. Ueno, A. Viets, L. Wade, M. Wade, and G. Waratkar, (2023), arXiv:2305.05625 [gr-qc] .EoS
- Sakon et al. (2022) S. Sakon et al., “Template bank for compact binary mergers in the fourth observing run of Advanced LIGO, Advanced Virgo, and KAGRA,” (2022), arXiv:2211.16674 [gr-qc] .EoS
- lig (2022) “Ligo sensitivity (190 mpc) (used for o4 simulations),” (2022).EoS
- Fong (2018) H. K. Y. Fong, From simulations to signals: Analyzing gravitational waves from compact binary coalescences, Ph.D. thesis, Toronto U. (2018).EoS
- Ray et al. (2023) A. Ray et al., “When to Point Your Telescopes: Gravitational Wave Trigger Classification for Real-Time Multi-Messenger Followup Observations,” (2023), arXiv:2306.07190 [gr-qc] .EoS
- Morisaki and Raymond (2020) S. Morisaki and V. Raymond, Phys. Rev. D 102, 104020 (2020), arXiv:2007.09108 [gr-qc] .EoS
- Cannon et al. (2010) K. Cannon, A. Chapman, C. Hanna, D. Keppel, A. C. Searle, and A. J. Weinstein, Phys. Rev. D 82, 044025 (2010).EoS
- Joshi et al. (2023) P. Joshi, L. Tsukada, and C. Hanna, “Background Filter: A method for removing signal contamination during significance estimation of a GstLAL anaysis,” (2023), arXiv:2305.18233 [gr-qc] .EoS
- Cannon et al. (2013) K. Cannon, C. Hanna, and D. Keppel, Phys. Rev. D 88, 024025 (2013), arXiv:1209.0718 [gr-qc] .EoS
- Joshi et al. (2025) P. Joshi, W. Niu, C. Hanna, R. Huxford, D. Singh, L. Tsukada, S. Adhicary, P. Baral, A. Baylor, K. Cannon, S. Caudill, M. W. Coughlin, B. Cousins, J. D. E. Creighton, B. Ewing, H. Fong, R. N. George, S. Ghosh, P. Godwin, R. Harada, Y.-J. Huang, C. Messick, S. Morisaki, D. Mukherjee, A. Pace, C. Posnansky, A. Ray, S. Sachdev, S. Sakon, U. Shah, R. Tapia, K. Ueno, A. Viets, L. Wade, M. Wade, Z. Yarbrough, and N. Zhang, “How many times should we matched filter gravitational wave data? a comparison of gstlal’s online and offline performance,” (2025), arXiv:2505.23959 [gr-qc] .EoS
- Chaudhary et al. (2023) S. S. Chaudhary et al., “Low-latency alert products and their performance in anticipation of the fourth ligo-virgo-kagra observing run,” (in prep) (2023).EoS
- Buonanno et al. (2009) A. Buonanno, B. R. Iyer, E. Ochsner, Y. Pan, and B. S. Sathyaprakash, Phys. Rev. D 80, 084043 (2009).EoS
- Bohé et al. (2017) A. Bohé, L. Shao, A. Taracchini, A. Buonanno, S. Babak, I. W. Harry, I. Hinder, S. Ossokine, M. Pürrer, V. Raymond, T. Chu, H. Fong, P. Kumar, H. P. Pfeiffer, M. Boyle, D. A. Hemberger, L. E. Kidder, G. Lovelace, M. A. Scheel, and B. Szilágyi, Phys. Rev. D 95, 044028 (2017).EoS
- Khan et al. (2016) S. Khan, S. Husa, M. Hannam, F. Ohme, M. Pürrer, X. J. Forteza, and A. Bohé, Phys. Rev. D 93, 044007 (2016).EoS
- Callister et al. (2017) T. Callister, J. Kanner, T. Massinger, S. Dhurandhar, and A. Weinstein, Classical and Quantum Gravity 34, 155007 (2017).EoS
- Ligo Scientific Collaboration et al. (2024a) Ligo Scientific Collaboration, VIRGO Collaboration, and Kagra Collaboration, General Coordinates Network 36236, 1 (2024a).EoS
- Note (1) The statistical significance of this candidate was subsequently estimated to be lower after collecting more noise triggers and the update alert was sent Ligo Scientific Collaboration et al. (2024b).EoS
- Messick (2019) C. Messick, Detecting Gravitational Waves for Multi-Messenger Astro, Ph.D. thesis (2019).EoS
- (62) L. S. Collaboration and V. Collaboration, “Gwtc-2 data release: Sensitivity of matched filter searches to binary black hole merger populations,” https://dcc.ligo.org/P2000217/public.EoS
- Abbott et al. (2022) R. Abbott, T. D. Abbott, F. Acernese, K. Ackley, C. Adams, N. Adhikari, R. X. Adhikari, V. B. Adya, C. Affeldt, D. Agarwal, et al., Astronomy & astrophysics 659, A84 (2022).EoS
- Ghosh et al. (2024) S. Ghosh, K. Chandra, and A. Pai, Physical Review D 109 (2024), 10.1103/physrevd.109.064015.EoS
- Macas et al. (2022) R. Macas, J. Pooley, L. K. Nuttall, D. Davis, M. J. Dyer, Y. Lecoeuche, J. D. Lyman, J. McIver, and K. Rink, Phys. Rev. D 105, 103021 (2022).EoS
- Ligo Scientific Collaboration et al. (2024b) Ligo Scientific Collaboration, VIRGO Collaboration, and Kagra Collaboration, General Coordinates Network 36812, 1 (2024b).EoS